OpenStudy (rainbow_dashie):

The product of 3 and a number is 12 more than the sum of the number and 26. What is the number?

Directrix (directrix):

Let x be the number. The product of 3 and a number: 3x sum of the number and 26: x + 26 12 more than the sum of the number and 26: x + 26 + 12 ----------------------------------------------- What is the equation we need to solve to get the value of x?

OpenStudy (rainbow_dashie):

@Directrix i'm not sure

Directrix (directrix):

Fill in for this. Use the variable expressions Use the symbols I posted that go with the words below. You can do it. (The product of 3 and a number) = (12 more than the sum of the number and 26}

OpenStudy (rainbow_dashie):

3x=? whats the rest?

OpenStudy (rainbow_dashie):

3x-12+26

OpenStudy (rainbow_dashie):

@jordanloveangel

OpenStudy (rainbow_dashie):

@StudyGurl14

OpenStudy (studygurl14):

3x = 12 more than the sum of the number and 26 3x = 12 + (x + 26) 3x = x + 26 + 12 Do the rest.

Directrix (directrix):

@Rainbow_dashie >>>3x=? whats the rest? The rest: 12 more than the sum of the number and 26: x + 26 + 12

OpenStudy (jordanloveangel):

3x=12+x+26 3x=x+38 2x=38 x=17

OpenStudy (jordanloveangel):

i think im not sure

OpenStudy (jordanloveangel):

To solve this problem, we must find where 3x - 12 = 26 + x Move all terms containing x to the left, all other terms to the right. Add '-1x' to each side of the equation. -12 + 3x + -1x = 26 + x + -1x Combine like terms: 3x + -1x = 2x -12 + 2x = 26 + x + -1x Combine like terms: x + -1x = 0 -12 + 2x = 26 + 0 -12 + 2x = 26 Add '12' to each side of the equation.

OpenStudy (jordanloveangel):

-12 + 12 + 2x = 26 + 12 Combine like terms: -12 + 12 = 0 0 + 2x = 26 + 12 2x = 26 + 12 Combine like terms: 26 + 12 = 38 2x = 38 Divide each side by '2'. x = 19 Simplifying x = 19 Check: 3*19 = 57 - 12 = 45 26 + 19 = 45